OpenStudy (anonymous):

Simplify 6* square root of 108 a. 54* square root of 6 b. 18* square root of 6 c. 36* square root of 3 d. 216* square root of 3

OpenStudy (anonymous):

@Posideon

OpenStudy (anonymous):

\[6*\sqrt{108}=6*\sqrt{9}*\sqrt{12}=6*3*\sqrt{12}=18*\sqrt{12}=18*\sqrt{4}*\sqrt{3}=18*2*\sqrt{3}=24\sqrt{3}\]

OpenStudy (anonymous):

my bad, last bit should say \[36\sqrt{3}\]